A solid metallic sphere of a radius a carries total charge Q.No other charges are nearby.The electric field just outside is surface is keQ/a squared radially outward. At this close point, the uniformly charged surface of the sphere looks exactly like a uniform flat sheet of charge. Is the electric field here given by o/e or o/2e?
1
Expert's answer
2020-03-24T09:57:01-0400

The electric field just outside its surface is 


"E=\frac{kQ}{a^2}"

"k=\frac{1}{4\pi \epsilon_0}"

"E=\frac{1}{ \epsilon_0}\frac{Q}{A}=\frac{\sigma}{ \epsilon_0}"